John cuts a cord of wood three times as fast as Ben. Together they can cut a cord of wood in 12 hours. How long does it take each one to cut a cord of wood?
Answer by jorel1380(3719)

1/t+3/t=1/12
12+36=t
t=48
Ben takes 48 hours to cut the wood. John takes 1/3rd of that, or 16 hours..